# Observation, Analysis, and Solution: Exploring Strong Lightweight Vision Transformers via Masked Image Modeling Pre-Training

### v1

@@ 5.3 Performance Improvements Based on Distillation
We finally evaluate our distillation-based MAE pre-trained models on all the classification and dense prediction tasks exploited in the observation section (Sec. 3). As shown in Tab. VI, our distillation contributes to significant performance improvements over the MAE pre-training without distillation, especially on the data-insufficient classification tasks and dense prediction tasks. In comparison to the supervised pre-training baseline (Our Recipe) relying on IN1K labels, our distillation-based MAE pre-training not only saves the costs for labeling the pre-training data, but also provides richer learning signals and avoids the bias towards the pre-defined set of IN1K labels. This leads to performance gains on Aircraft, Cars, iNat18, etc., which may contain less overlapped categories or domains with ImageNet than Flowers, Pets and CIFAR100. The decoupled distillation strategy D2-MAE also consistently improves over D-MAE on all the evaluated datasets.
@@ 6 State-of-the-art Comparison Results
Our observation and analysis on Hiera-Tiny are almost the same as that on ViT-Tiny, so we can directly apply our decoupled distillation on its MAE pre-training to improve the performance. It is noteworthy that we evaluate performance for Hiera-Tiny in this section only on the ImageNet classification task and COCO object detection and instance segmentation tasks following the original paper to achieve the image results due to its intrinsic hierarchical property. In Sec. C.1, we also show that our distillation technique D-MAE can help other-scale vanilla ViT students beyond ViT-Tiny to achieve better downstream performance.
TABLE VII: Comparison Results to the Current SOTA Lightweight Models on ImageNet-1K. Top-1 Accuracy on Its Validation Set with Parameter Count and Latency Are Reported. All Models Are Tested on Input Scale of 224$\times$224, Except That MobileViT [21] and FastViT [24] Are Tested with 256$\times$256 According to Their Original Papers. ‘bsn’ Means the Latency Is Measured with Batch Size n.
@@
Methods Data Distill. #Epochs Orin (bs1) A100 (bs16) (%)
@@
TABLE VIII: Comparison Results to Previous SOTA Lightweight Models in the Mask R-CNN Framework on COCO.
@@
ResNet18 [74] 31M 207G Super. 37.1 57.3 40.0 34.0 54.3 36.4
ResNet50 [74] 44M 260G Super. 41.0 61.7 44.9 37.1 58.4 40.1
PVT-T [88] 33M 208G Super. 39.8 62.2 43.0 37.4 59.3 39.9
LightViT-T [146] 28M 187G Super. 41.5 64.4 45.1 38.4 61.2 40.8
Super. 40.4 60.7 43.1 35.5 56.8 36.9
ViT-Tiny 28M 204G D2-MAE 42.5 62.6 45.6 37.5 59.4 39.4
Super. 37.9 58.7 40.4 34.3 55.5 35.9
Hiera-Tiny 28M 205G D2-MAE 43.2 64.0 46.7 38.7 60.7 41.4
TABLE IX: Comparison Results to Lightweight Semantic Segmentation Methods on ADE20K.
Backbones #Param. FLOPs Pre-Train mIoU
High-resolution Segmentation-specific
SeaFormer-L [27] 14.0M 6.5G Super. 42.7
EfficientViT-B1 [26] 4.8M 6.3G Super. 42.8
Versatile Backbones with Semantic-FPN Decoder
ResNet18 [74] 15.4M 32.3G Super. 34.1
LeViT-192 [17] 14.7M 10.2G Super. 34.8
PVT-T [88] 17M 16.6G Super. 35.7
LightViT-T [146] 12.2M 33.1G Super. 39.9
Super. 38.0
D2-MAE 39.0
Super. 41.5
ViT-Tiny 7.0M 13.5G D2-MAE 42.8
TABLE X: Comparison Results to Previous SOTA Lightweight CPU-Realtime Trackers on LaSOT. We Report CPU Runtime Speeds in FPS.
Trackers Backbones #Param. FLOPs Pre-Train AUC Speed
HiT-Base [147] LeViT-384 42.1M 4.3G Super. 64.6 33
HiT-Small [147] LeViT-128 11.0M 1.1G Super. 60.5 72
HiT-Tiny [147] LeViT-128S 9.6M 1.0G Super. 54.8 76
LightTrack-L [148] LT-Mobile 3.1M 0.8G Super. 55.5 -
LightTrack-M [148] LT-Mobile 2.0M 0.5G Super. 53.8 41
E.T.Track [149] LT-Mobile 7.0M 1.7G Super. 59.1 47
FEAR-L [150] RegNet 4.1M 4.9G Super. 57.9 -
FEAR-XS [150] FBNet 1.0M 1.0G Super. 53.5 60
HCAT [151] ResNet18 6.8M 1.3G Super. 59.3 45
MixFormerV2-S [152] ViT-Base 16.2M 4.4G Super. 60.6 30
Super. 64.1 41
D-MAE 65.8 41
OSTrack-T ViT-Tiny 8.0M 1.9G D2-MAE 66.1 41
  • $\bm{\star}$
@@
    indicates the models adopt an intermediate 1000-epoch fine-tuning on IN1K classification before ADE20K segmentation fine-tuning.
6.1 Comparison on ImageNet Classification
@@
In Tab. VII, we compare both the enhanced ViT-Tiny and Hiera-Tiny with our distillation-based MAE pre-training to the current lightweight ConvNets and ViT derivatives. We report top-1 accuracy along with the model parameter count and the latency. By taking inspiration from the commonly used strategies in the network architecture design methodology, such as extending the supervised fine-tuning duration to a much longer schedule and using relative position bias, our approach can further benefit from them to achieve $79.4\%$/$78.9\%$ top-1 accuracy on ImageNet-1K, which is comparable to the current SOTA networks with sophisticated architecture design. It means that our proposed distillation strategy for the MAE pre-training can unleash the great potential of the extremely simple lightweight ViTs, which is commonly overlooked in the past several years. Although some of the hierarchical ViTs (e.g., LeViT-192, EfficientFormerV2-S1 and EfficientViT-B1) enjoy higher accuracy and more attractive latency than ours simultaneously, our pre-training enables the more versatile ViT model ViT-Tiny to be outstanding on both the ADE20K semantic segmentation and LaSOT single object tracking tasks in the lightweight regime (see Sec. 6.3 and Sec. 6.4 below).
6.2 Comparison on COCO Detection and Segmentation
@@
For a more thorough study, we further conduct the SOTA comparison on the downstream object detection and segmentation tasks of COCO [132]. As we have detailed in Sec. 3.3, we reproduce and rigorously follow the benchmarking setup in the Mask R-CNN framework in [135] except for some minimal changes to the training settings. In specific, the different models in Tab. VIII are used for the initialization of the backbone. Our pre-training not only improves over the supervised pre-training significantly, but also makes the extremely simple hierarchical Hiera-Tiny to outperform the delicately designed convolution-free ViT derivatives PVT-T and LightViT-T.
6.3 Comparison on ADE20K Semantic Segmentation
@@
We also conduct the SOTA comparison on the downstream semantic segmentation task of ADE20K [133]. To align with the recent effort on reducing the computational cost of high-resolution dense prediction to ease its deploying on hardware devices, e.g., the segmentation-specific models SeaFormer and EfficientViT, we instead use the Semantic-FPN framework [153] to evaluate the different lightweight versatile backbones in Sec. 6. It can be seen that the image classification-proficient model LeViT with decreasing resolutions can not perform well on this ADE20K dense prediction task, while our pre-training with intermediate fine-tuning on IN1K enables ViT-Tiny to achieve comparable accuracy with the segmentation-specific models.
6.4 Comparison on LaSOT Visual Tracking
@@
We finally conduct the SOTA comparison on the downstream single object visual tracking task of LaSOT [133] in the lightweight regime, in which the real-time running on limited computing resources such as CPU is required. As we have detailed in Sec. 3.3, we reproduce and rigorously follow the original setup in the speed-oriented version of the SOTA tracker OSTrack [137] (256$\times$256 input size of search patch) except that we only replace the backbone for joint feature extraction and relation modeling with different pre-trained ViT-Tiny, namely OSTrack-T in Sec. 6. It shows that our pre-training enhances OSTrack-T to outperform all the current SOTA lightweight CPU-realtime trackers and set a new record of $66.1\%$ AUC on LaSOT.
@@ 7 Conclusion
In this paper, we investigate an easily overlooked aspect of obtaining strong lightweight ViTs through MIM pre-training, which is optimizing the pre-training strategies for improving the extremely simple lightweight ViTs and hence orthogonal to the delicate design of ViT derivatives by introducing sophisticated components with inductive biases. Our comprehensive analyses and investigations reveal several key factors that hide behind the achieved superior results, the value of which has also been demonstrated as it guides the design of our proposed distillation strategy for better MIM pre-training. We believe the observations, analyses and solutions in this paper will improve the understandings of MIM pre-training for lightweight ViTs, and we expect our observation-analysis-solution flow will further advance the SSL-based pre-training research.
Supplementary Material
This is a supplementary material for our paper “Observation, Analysis, and Solution: Exploring Strong Lightweight Vision Transformers via Masked Image Modeling Pre-Training”.
@@ A.2 Transfer Evaluation Details for Classification Tasks
Appendix B More Analyses on The Pre-Training
@@ B.1 Analyses with More Models as Reference
In the main paper, the analyses based on the layer-wise representation similarity are mainly conducted by adopting the fully-supervised ViT-Tiny trained on the IN1K classification task based on our recipe (see Tab. I) as the reference model. Here, we additionally introduce stronger recognition models as references to demonstrate the generalizability of our analyses. Specifically, we use ViT-Base models trained with various recipes as references, e.g., DeiT-Base (supervised training on IN1K following [16] which achieves 82.0% top-1 accuracy on IN1K), ViT-Base-21k (supervised pre-training on IN21K following [70]), ViT-Base-21k-1k (first supervised pre-training on IN21K and then fine-tuning on IN1K following [70], achieving 84.5% top-1 accuracy on IN1K). The layer representation similarity is presented in Fig. S1.
@@ C.1 Applying Distillation on More Networks
Models Channel Dimension #Heads #Layers #Params

### v2

@@ 5.3 Performance Improvements Based on Distillation
We finally evaluate our distillation-based MAE pre-trained models on all the classification and dense prediction tasks exploited in the observation section (Sec. 3). As shown in Tab. 6, our distillation contributes to significant performance improvements over the MAE pre-training without distillation, especially on the data-insufficient classification tasks and dense prediction tasks. In comparison to the supervised pre-training baseline (Our Recipe) relying on IN1K labels, our distillation-based MAE pre-training not only saves the costs for labeling the pre-training data, but also provides richer learning signals and avoids the bias towards the pre-defined set of IN1K labels. This leads to performance gains on Aircraft, Cars, iNat18, etc., which may contain less overlapped categories or domains with ImageNet than Flowers, Pets and CIFAR100. The decoupled distillation strategy D2-MAE also consistently improves over D-MAE on all the evaluated datasets.
@@ 6 State-of-the-art Comparison Results
Table 7: Comparison results to the current SOTA lightweight models on ImageNet-1K. Top-1 accuracy on its validation set with parameter count and latency are reported. All models are tested on input scale of 224$\times$224, except that MobileViT [21] and FastViT [24] are tested with 256$\times$256 according to their original papers. ‘bsn’ means the latency is measured with batch Size n.
@@
Methods Data Distill. Epochs Orin(bs1) A100(bs16) (%)
@@
Our observation and analysis on Hiera-Tiny are almost the same as that on ViT-Tiny, so we can directly apply our decoupled distillation on its MAE pre-training to improve the performance. It is noteworthy that we evaluate performance for Hiera-Tiny in this section only on the ImageNet classification task and COCO object detection and instance segmentation tasks following the original paper to achieve the image results due to its intrinsic hierarchical property. In Appendix C.1, we also show that our distillation technique D-MAE can help other-scale vanilla ViT students beyond ViT-Tiny to achieve better downstream performance.
@@
[…]
[rotAngle=90,nonFloat=true,capWidth=w,capPos=top]table Backbones #Param. FLOPs Pre-Train mIoU High-resolution Segmentation-specific SeaFormer-L [27] 14.0M 6.5G Super. 42.7 EfficientViT-B1 [26] 4.8M 6.3G Super. 42.8 Versatile Backbones with Semantic-FPN Decoder ResNet18 [79] 15.4M 32.3G Super. 34.1 LeViT-192 [17] 14.7M 10.2G Super. 34.8 PVT-T [93] 17M 16.6G Super. 35.7 LightViT-T [146] 12.2M 33.1G Super. 39.9 Super. 38.0 D2-MAE 39.0 Super. 41.5 ViT-Tiny 7.0M 13.5G D2-MAE 42.8 Comparison results to lightweight semantic segmentation methods on ADE20K.tab:ade20ksota
[…]
[rotAngle=90,nonFloat=true,capWidth=w,capPos=top]table Trackers Backbones #Param. FLOPs Pre-Train AUC Speed HiT-Base [147] LeViT-384 42.1M 4.3G Super. 64.6 33 HiT-Small [147] LeViT-128 11.0M 1.1G Super. 60.5 72 HiT-Tiny [147] LeViT-128S 9.6M 1.0G Super. 54.8 76 LightTrack-L [148] LT-Mobile 3.1M 0.8G Super. 55.5 - LightTrack-M [148] LT-Mobile 2.0M 0.5G Super. 53.8 41 E.T.Track [149] LT-Mobile 7.0M 1.7G Super. 59.1 47 FEAR-L [150] RegNet 4.1M 4.9G Super. 57.9 - FEAR-XS [150] FBNet 1.0M 1.0G Super. 53.5 60 HCAT [151] ResNet18 6.8M 1.3G Super. 59.3 45 MixFormerV2-S [152] ViT-Base 16.2M 4.4G Super. 60.6 30 Super. 64.1 41 D-MAE 65.8 41 OSTrack-T ViT-Tiny 8.0M 1.9G D2-MAE 66.1 41 Comparison results to previous SOTA lightweight CPU-realtime trackers on LaSOT. We report CPU runtime speeds in FPS.tab:lasotsota
[…]
@@
[rotAngle=90,nonFloat=true,onlyText=true] table $\bm{\star}$ indicates the models adopt an intermediate 1000-epoch fine-tuning on IN1K classification before ADE20K segmentation fine-tuning. fig:resqrbd
@@
Comparison on ImageNet Classification. In Tab. 7, we compare both the enhanced ViT-Tiny and Hiera-Tiny with our distillation-based MAE pre-training to the current lightweight ConvNets and ViT derivatives. We report top-1 accuracy along with the model parameter count and the latency. By taking inspiration from the commonly used strategies in the network architecture design methodology, such as extending the supervised fine-tuning duration to a much longer schedule and using relative position bias, our approach can further benefit from them to achieve $79.4\%$/$78.9\%$ top-1 accuracy on ImageNet-1K, which is comparable to the current SOTA networks with sophisticated architecture design. It means that our proposed distillation strategy for the MAE pre-training can unleash the great potential of the extremely simple lightweight ViTs, which is commonly overlooked in the past several years. Although some of the hierarchical ViTs (e.g., LeViT-192, EfficientFormerV2-S1 and EfficientViT-B1) enjoy higher accuracy and more attractive latency than ours simultaneously, our pre-training enables the more versatile ViT model ViT-Tiny to be outstanding on both the ADE20K segmentation and LaSOT tracking tasks in the lightweight regime (see below).
@@
Comparison on COCO Detection and Segmentation. For a more thorough study, we further conduct the SOTA comparison on the downstream object detection and segmentation tasks of COCO [34]. As we have detailed in Sec. 3.3, we reproduce and rigorously follow the benchmarking setup in the Mask R-CNN framework in [136] except for some minimal changes to the training settings. In specific, the different models in LABEL:tab:cocosota are used for the initialization of the backbone. Our pre-training not only improves over the supervised pre-training significantly, but also makes the extremely simple hierarchical Hiera-Tiny to outperform the delicately designed convolution-free ViT derivatives PVT-T and LightViT-T.
@@
Comparison on ADE20K Semantic Segmentation. We also conduct the SOTA comparison on the downstream semantic segmentation task of ADE20K [35]. To align with the recent effort on reducing the computational cost of high-resolution dense prediction to ease its deploying on hardware devices, e.g., the segmentation-specific models SeaFormer and EfficientViT, we instead use the Semantic-FPN framework [153] to evaluate the different lightweight versatile backbones in LABEL:tab:ade20ksota. It shows that the image classification-proficient model LeViT with decreasing resolutions can not perform well on this ADE20K dense prediction task, while our pre-training with intermediate fine-tuning on IN1K enables ViT-Tiny to achieve comparable accuracy with the segmentation-specific models.
@@
Comparison on LaSOT Visual Tracking. We finally conduct the SOTA comparison on the downstream single object visual tracking task of LaSOT [35] in the lightweight regime, in which the real-time running on limited computing resources such as CPU is required. As we have detailed in Sec. 3.3, we reproduce and rigorously follow the original setup in the speed-oriented version of the SOTA tracker OSTrack [37] (256$\times$256 input size of search patch) except that we only replace the backbone for joint feature extraction and relation modeling with different pre-trained ViT-Tiny, namely OSTrack-T in LABEL:tab:lasotsota. It shows that our pre-training enhances OSTrack-T to outperform all the current SOTA lightweight CPU-realtime trackers and set a new record of $66.1\%$ AUC on LaSOT.
@@ 7 Conclusion
In this paper, we investigate an easily overlooked aspect of obtaining strong lightweight ViTs through MIM pre-training, which is optimizing the pre-training strategies for improving the extremely simple lightweight ViTs and hence orthogonal to the delicate design of ViT derivatives by introducing sophisticated components with inductive biases. Our comprehensive analyses and investigations reveal several key factors that hide behind the achieved superior results, the value of which has also been demonstrated as it guides the design of our distillation strategy for better MIM pre-training. We believe our observations, analyses and solutions will improve the understanding of MIM pre-training for lightweight ViTs. We expect our observation-analysis-solution flow will further advance the SSL-based pre-training research.
@@ A.2 Transfer Evaluation Details for Classification Tasks
Appendix B More Analyses on Pre-Training
@@ B.1 Analyses with More Models as Reference
In the main paper, the analyses based on the layer-wise representation similarity are mainly conducted by adopting the fully-supervised ViT-Tiny trained on the IN1K classification task based on our recipe (see Tab. 1) as the reference model. Here, we additionally introduce stronger recognition models as references to demonstrate the generalizability of our analyses. Specifically, we use ViT-Base models trained with various recipes as references, e.g., DeiT-Base (supervised training on IN1K following [16] which achieves 82.0% top-1 accuracy on IN1K), ViT-Base-21k (supervised pre-training on IN21K following [75]), ViT-Base-21k-1k (first supervised pre-training on IN21K and then fine-tuning on IN1K following [75], achieving 84.5% top-1 accuracy on IN1K). The layer representation similarity is presented in Fig. S1.
@@ C.1 Applying Distillation on More Networks
Models Channels #Heads #Layers #Params